OpenStudy (yamyam70):

Factor out the GCF 5x^4 (x-y) -3x^3y^5(y-x) they said something about the "changing sign" I do not know where to start . help anyone ?

OpenStudy (dan815):

y-x = -*(x-y)

OpenStudy (yamyam70):

did you do that to the whole given?

OpenStudy (yamyam70):

I mean do you apply that to the whole

OpenStudy (dan815):

=5x^4 (x-y) +3x^3y^5(x-y) =x^3*(x-y)(5x+3y^5)

OpenStudy (yamyam70):

did you just multiply the negative beside the 3 to the binomial ?

OpenStudy (yamyam70):

Thank you so much :)